(SE 2ГЕР — - / k/


Cantlem Exemple: isky 25kp B с fe a) and support Reaction 1) Draw Sf.D () Draw B.mo
Answer #1

Ans :

Resultant reaction forces acting at support is given below :

ADS: FB D for Beam: 15 kipf 25 kipf A Xc 714 - 1 4ft 4 ft 4 Ft TRAV 7 RAH S MA * for cantilever beam, at fixed support there

1) To find reaction forces at bupport consider Efx = 0 =tve =-ve). RAH=0 Now p=tve Efx=0 ** x=-ve 9 RAV= 15+25 RAV=40 kipf RA

SFD and BMD calculations and graph is given below :

93 for eg shear force dia. calculations. There are different methods of solving we will make 2 compartments around force left

* SFD calculations (1=tve, v=-ve SFAL= ARE 40 kipf. BL = 490 kipf BR = yo-15 25 kipf. CLE 25 kipf CR = 25 -25 = o kipf o 11 k

?! BM А 2508 - ) 5x4 x BIX) - 260 kipf.ft Now BMB 25X4 -ioo BMB * .ioo kipfo ft There is no force right t to ptoc BMc=0 es

SFD and BMD :

P = 15 P = 25 НА A МА 4 (ft) + 4 (ft) + 4 (ft) * * 1 sec. (X1), 2 sec. (X2) 3 sec. (X3) 40 25 Оy 0 -260 -100 Mx